OpenStudy (anonymous):

f(x)= 2x^2+2x+10 find the vertex and line of symmetry

OpenStudy (dumbcow):

\[x = \frac{-b}{2a}\] this gives line of symmetry and x_coord of vertex

OpenStudy (anonymous):

would x be 1/2

OpenStudy (anonymous):

dumbcow is correct on the x coord. For the y coord it is the f(-b/2a)

OpenStudy (anonymous):

x=-1/2

OpenStudy (anonymous):

\[y=2(x^2+x+5)\]

OpenStudy (anonymous):

x= 1/2 which would also be the line of symetry

OpenStudy (anonymous):

\[y=2(x+\frac{1}{2})^2+5-(2*\frac{1}{4})\]

OpenStudy (anonymous):

-1/2 not +1/2 and plug that back into the original equation to get Y then you will have (-1/2, Y) = Vertex

OpenStudy (anonymous):

in the end y should be an improper fraction

OpenStudy (anonymous):

2x^2+2x+10 = 2(-1/2)^2 + 2(-1/2) + 10 = y

OpenStudy (anonymous):

i get -3/2 is that correct?

OpenStudy (anonymous):

for the y

OpenStudy (dumbcow):

no y should be 9.5 2(-1/2)^2 + 2(-1/2) + 10

OpenStudy (anonymous):

it should be in simplified fraction form

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2%28-1%2F2%29^2+%2B+2%28-1%2F2%29+%2B+10 according to that you shouldve got 19/2

OpenStudy (anonymous):

it said the correct answer was 21/2

OpenStudy (anonymous):

The book is on drugs

OpenStudy (anonymous):

\[y=2(x+\frac{1}{2})^2+\frac{21}{2}\] \[vertex: ( -\frac{1}{2}, \frac{21}{2})\] panther my answer is correc with your key

OpenStudy (anonymous):

i only have the answer in the book to go from

OpenStudy (anonymous):

look at my other post and tell me if you agree with my answers

OpenStudy (anonymous):

x= -1/2 y= 21/2

OpenStudy (anonymous):

ok ,i will try